OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (anonymous):

I would set up the 3x4 matrix and reduce it to reduced row echelon form to solve the system of equation. Now that I've answered that question for you, do you have any questions about that question?

OpenStudy (anonymous):

Do you have time to do this step by step?

OpenStudy (anonymous):

The method I used is matrix algebra. Have you learned how to do that yet?

OpenStudy (anonymous):

Isn't that where you have to use two eliminate two equations solve for the variable then use the answer that i get from that to plug it into the equation that i didn't use in the beginning?

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -5 & 2 & 0 \\ 1 & 4 & -1 & 12\\ 2 & -1 & 3 & 10\end{matrix}\right]\] is the matrix that you want to set up. Reduced row echelon form is \[\left[\begin{matrix}1 & 0 & 0 & a \\ 0 & 1 & 0 & b\\ 0 & 0 & 1 & c\end{matrix}\right]\] where \[\left[\begin{matrix}x & y & z & constant \end{matrix}\right]\] so the solution is x =a , y =b z= c

OpenStudy (anonymous):

so they all = 1?

OpenStudy (anonymous):

you reduce it to 1. You take the second row, and subtract it from the first (1 4 -1 12) - (1 -5 2 0) and you get (0 9 -3 12) You take the third row, and subtract it from two times the first row (2 -1 3 10) - 2*(1 -5 2 0) and you get 0 9 -1 0) and now you're left with this matrix \[\left[\begin{matrix}1 & -5 & 2 & 0 \\ 0 & 9 & -3 & 12\\ 0 & 9 & -1 & 10\end{matrix}\right]\] the subtract the third row from the second row \[\left[\begin{matrix}1 & -5 & 2 & 0 \\ 0 & 9 & -3 & 12\\ 0 & 0 & 2 & -2\end{matrix}\right]\] Divide the third row by 2 \[\left[\begin{matrix}1 & -5 & 2 & 0 \\ 0 & 9 & -3 & 12\\ 0 & 0 & 1 & -1\end{matrix}\right]\] and now you have a simpler system of equations x - 5y +2z = 0 0x+9y -3z = 12 0x + 0y +z = -1 Then substitute the rest or continue manipulating the matrix

OpenStudy (anonymous):

ooh okay thanks :)

OpenStudy (radar):

Have you had that method yet @Yoduh ?

OpenStudy (anonymous):

To be honest i haven't but thing is i want to do this using the Elimination method. Think you can help me?

OpenStudy (radar):

The elimination method is certaInly a method that could be used. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (radar):

Note that the first two equations can be used to eliminate x. as you can do this x=x or 5y-2z=-4y + z + 12 or 9y - 3z - 12 = 0 see the x is gone now we need to eliminate y or z. Did you see how the x was eliminated?

OpenStudy (anonymous):

yes i know how to do that part just it gets a little confusing when i want to eliminate whichever other variable. I get two variables on one side and a regular number on the other

OpenStudy (anonymous):

Give me one second I'm going to do this in equation form so you can see if I'm doing it right

OpenStudy (anonymous):

x-5y+2z=0 and x+4y-z=12 I need to eliminate the x so want to multiply which ever equation by -1 -1(x-5y+2z=0) making it -x+5y-2z=0 then i add them -x+5y-2z=0 + x+4y-z=12 9y-z=12 am I correct?

OpenStudy (anonymous):

or is t 9y-3z=12

OpenStudy (radar):

the latter

OpenStudy (radar):

addition of (-2z and -z)= -3z

OpenStudy (radar):

You have now eliminated x and have the equation 9y-3z=12

OpenStudy (anonymous):

great but this is what i mean two variables on one side? how do i fix that into making it to one variable so i can use it as part of the equation?

OpenStudy (radar):

As a suggestion take the first and last equation and eliminate x again giving you another new equation with only y and z terms.

OpenStudy (anonymous):

i got this X+5y+2z=0 +2x-y+3z=10 = 3x-4y+5z=10

OpenStudy (radar):